OpenStudy (anonymous):

A rectangular display case houses a square pyramid. Both have the same square base measuring x inches on a side. The top of the case is 2 inches above the tip of the pyramid which has a height of h inches. Which expression represents the volume of the empty space inside the case? Answer: a. x^3 - x^2(h-2) b. x^2-1/3x^2h c. 1/3x2^h d. 2/3x^2h e. x^2(h+2) - 1/3x^2h

OpenStudy (mayankdevnani):

The rectangular display case has volume (x)(x)(h+2) = x^2h + 2 x^2. The housed pyramid has volume 1/3 B h = 1/3 (x*x) h= 1/3x^2h. Empy Space = Case volume - pyramid volume = x^2 h + 2 x^2 - 1/3 x^2 h = x^2 h - 1/3 x^2 h + 2 x ^2 = [ (2/3) x^2 h + 2 x ^ 2] cubic inches

OpenStudy (mayankdevnani):

right @thivitaa

OpenStudy (mayankdevnani):

Step 1) Find the volume of a rectangular prism Step 2) Find the volume of a pyramid with square base Setp 3) (rectangular prism volume) MINUS (pyramid with square base volume).

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

im writing notes